OpenStudy (anonymous):

Which table shows a proportional relationship between x and y? 1. x: 1 3 4 6 y: 5 12 18 30 2. x: 1 2 4 5 y: 3 6 12 15 3. x: 0 2 4 7 y: 4 8 12 18 4. x: 1 2 3 6 y: 4 10 12 26 I think it's the third one but can someone help me just to be sure?

OpenStudy (anonymous):

WILL MEDAL!!

OpenStudy (tylerd):

if someonthing is proportional it looks like this y=kx

OpenStudy (tylerd):

so see if you can come up with an equation in the form y=kx for one of those tables

OpenStudy (anonymous):

what does the k stand for?

OpenStudy (tylerd):

its some constant, could be any number

OpenStudy (tylerd):

try to see if you can solve for k, for #2

OpenStudy (anonymous):

I..... can't. I'm super sorry :(

OpenStudy (anonymous):

I'm not the best in math

OpenStudy (tylerd):

so lets look at the numbers for #2. when x=1 y=3 when x=2 y=6 when x=4 y=12 so for y=kx we substitute the values for x and y in 3=k*1 6=k*2 12=k*4

OpenStudy (tylerd):

if k is the same value for each of those, then its proportional

OpenStudy (anonymous):

I think it is a proportional relationship

OpenStudy (anonymous):

Is it @TylerD ?